# Intrinsic back-switching phenomenon in SOT-MRAM devices

### v2

@@ I Introduction
Magnetic random-access memories (MRAMs) are considered one of the most promising replacements for volatile CMOS memory technologies across all levels of the memory hierarchy. Among the latest generations of MRAMs, spin-orbit torque MRAMs (SOT-MRAMs) are particularly attractive for replacing memories closest to the processing units, as they combine high endurance with sub-nanosecond switching times [1, 2]. Nevertheless, while spin transfer torque MRAMs (STT-MRAMs) are commercially available [3], SOT-MRAMs must overcome several challenges before they can be adopted by the industry [2]. One of these major challenges is to achieve deterministic switching, meaning reliably controlling the result of the write operation, over a wide current range. The key metric for deterministic switching is the write error rate, which measures the probability of getting the expected final state after a write operation.
@@
In SOT-MRAM technology with perpendicular magnetic anisotropy, which is the most promising solution in terms of scaling, density and data retention, magnetization switching by means of spin-orbit current alone is not deterministic. Indeed, by symmetry, the probabilities of obtaining an ”up” or ”down” magnetization are equal. An additional factor to break this symmetry is required. In early implementations, this breaking of symmetry was induced by the application of an in-plane magnetic field, collinear with the current but external to the system [4, 5, 6]. As this solution is not suitable for embedded applications, a number of propositions have been published in recent years, aimed at developing deterministic and integrated ”field-free” SOT-MRAM cells, based on tilted anisotropy [7], combination of SOT and STT effects [8], exchange bias [9], embedded magnet in the material stack [10] and unconventional torques from new materials [11] (for a more complete list, please consult review articles such as [12]).
@@
However, the presence of symmetry breaking does not guarantee a perfectly deterministic reversal of magnetization, and consequently a deterministic writing. It has been observed that increasing the write current beyond the threshold for reversal leads, as expected, to an increase in the probability of magnetization switching. However, if the current is increased further, the probability then decreases and may even oscillate as a function of the applied current [13, 14, 15]. Similar back-switching (BSW) phenomenon has previously been observed and studied in STT-MRAMs [16]. One origin of this phenomenon has been attributed to the instability of the reference layer magnetization in magnetic tunnel junctions (MTJ), an instability that can even lead to the reversal of this magnetization [17, 18]. In SOT-MRAMs, magnetization is switched by means of a current injected in the plane of the layers [4, 6], without the need for an MTJ-type structure to obtain a spin-polarized current. Observation of the back-switching phenomenon in simple structures, such as non-magnetic/ferromagnetic/insulator trilayer stacks, demonstrates that in SOT-MRAM this BSW is intrinsic to the magnetization dynamics induced by spin-orbit torques and not related to the dynamics of the reference layer. This BSW phenomenon is a challenge for the efficient design of memories based on SOT-MRAMs (Fig. 1 (a)). Its study is essential to our understanding of the magnetization dynamics induced by spin-orbit torques, and for finding effective means of mitigating BSW in SOT-MRAMs.
@@
Despite its importance, relatively few studies have been devoted to it. Most are numerical studies based on macrospin simulations. The existence of this BSW phenomenon was reported as early as the first macrospin SOT switching simulations, whether with or without the Field-Like (FL) term [19, 20]. The importance and the role of the FL term on the BSW phenomenon were then highlighted [21, 22, 23]. The existence of this BSW phenomenon has been explained by energy arguments: at the end of the write current pulse, if the energy of the system is higher than the saddle point energy, the magnetization will relax in either a first potential well corresponding to a first direction of magnetization or in the other corresponding to the opposite state of magnetization [20, 22]. This explains the high sensitivity of the phenomenon to damping, as reported in particular by [19, 20, 24]. Experimentally, this phenomenon has been reported in some publications - for example, in Pt/Co/AlO layers, where the magnetization is reported to switch back to its original orientation above a certain threshold of applied planar magnetic field [15]; or in Ta/CoFeB/MgO-based magnetic tunnel junctions, where a decrease in the probability of the magnetization switching is observed when the write current is increased [25]. Only two publications from the same group report a more complete study of this phenomenon, combining experiments and simulations [14, 13]. Both studies were carried out on micron-sized devices patterned from a Ta/CoFeB/MgO based stack. Time-resolved MOKE measurements show a decrease in the probability of switching as the writing current increases. Using analytical modeling and micromagnetic simulations, these measurements were explained by a BSW phenomenon due to the reflection of the magnetic domain walls (DW) on the edges of the device. This reflection can only occur for a negative sign of the ratio of the FL and Damping-Like (DL) torques [14]. In a subsequent study, based on electrical measurements on Hall crosses supporting a magnetic pillar, the authors observed an oscillatory and deterministic behavior of the magnetization switching while increasing the amplitude of nanosecond write pulses [13].
@@
In this work, we present a systematic study of the back-switching phenomenon using a statistical approach. The measurements were carried out on sub-100nm CoFeB magnetic pilars at the center of -W Hall crosses. The error rate of the write operation was measured as a function of various parameters such as the amplitude of the write current pulse and an external magnetic field applied along different directions. We show that macrospin simulations, using realistic values for the different magnetic parameters, reproduce the experimental observations well. These macrospin simulations open the door to the development of compact models that take this BSW phenomenon into account. These numerical simulations give us an insight into the mechanism behind the BSW phenomenon. They also allow us to study ways of reducing this phenomenon, such as using stronger damping, tuning the FL/DL ratio or, following previous work from [23] controlling the shape of the writing pulses. We test the latter solution and show experimentally that this strategy does indeed mitigate back-switching and increases the current range for write operations in SOT-MRAMs.
@@ II Methodology
The measurements were carried out on a stack constituting the typical storage layer of a SOT-MRAM: -W(4)/CoFeB(0.9)/MgO(RA20)/CoFeB(0.3)/Ta(1)/ Ru(5) where the thicknesses are expressed in nanometers. The top 0.3 nm CoFeB layer is non-magnetic and is used to provide appropriate environment to the MgO layer during annealing. The layers were sputter deposited and annealed at 350° C for 30 minutes unless otherwise specified in the text. This stack was then patterned as a CoFeB/MgO/CoFeB/Ta/Ru pillar at the center of a -W Hall cross by successive steps of electron lithography and ion beam etching. The studies reported here, except for certain cases specified in the text, were carried out on pillars with a real diameter of around 60 nm siting on a 100 nm wide track. Fig. 1 (b) presents the magnetic stack, a representative SEM image of a studied device, and a schematic of the measurement setup.
The study is based on statistical measurements of Write Error Rate (WER), which is a crucial metric for qualifying memory devices. Typically, while this WER is very low in SRAM memories, it is much higher in MRAMs. However, to operate as a working memory, the WER of a single bit should typically be less than 10-9 if the chip incorporates error correction code (ECC), or less than 10-18 without this error correction [26]. Fig. 2 (a) shows the WER measurement scheme that has been used and its definition: the memory point is initialized to a defined state and this state is read to verify that initialization was successful. A pulse to write the opposite state, of variable amplitude, is then applied and the state of the memory point is read again. WER is defined as the ratio of the number of write errors after a successful initialization to the number of successful initializations. Note that a WER of 0.5 corresponds to an equal probability of obtaining the two states of opposite magnetization.
@@ III Experimental Results
Magnetic characterization of the samples confirms perpendicular anisotropy, as evidenced by square hysteresis loops, shown in [27]. The coercive field is approximately 750 Oe, while the anisotropy field (H) and the thermal stability factor (), extracted using the switching field distribution method [28, 29], are estimated to be around H=2600 Oe and =45, respectively [27]. A typical hysteresis curve for current induced magnetization reversal is shown in Fig. 2 (b). The areas colored in green represent known areas of magnetization reversal into the opposite state. However, as the amplitude of the applied voltage increases, a new behavior appears, represented by the areas colored in red. The final state of magnetization is no longer fixed, opposite to the initial state, but fluctuates in an apparently random manner between these two states.
@@
The corresponding WER measurement is shown in Fig. 2 (c). This curve was obtained with 100 write/read cycles, which explains the saturation at WER, representing in this case an absence of error. Starting from a down state and applying a positive voltage, the WER of the up state is measured. We first observe a zone without switching when the voltage is below a first threshold V0. Then when the voltage is increased above V0, we observe that, after a transition zone, the magnetization switches to the up state without error, corresponding to the green zone in Fig. 2 (b). When the voltage increases further, there is another threshold V1 at which the error rate starts increasing and approaches the maximum value 1, which corresponds to the magnetization switching back to its initial orientation, in this case down, without any error. This behavior is repeated when the magnetization is initialized in the opposite up state and by measuring the WER of the down state. By applying a positive voltage below V1, the up state remains stable while above V1, the magnetization switches to the down state. This shows that the BSW is not random but, on the contrary, deterministic: after being reversed to the orientation opposing the initial state, the magnetization can be switched back to its initial orientation in a deterministic way, as shown by the low error rate obtained (cf Fig. 2 (c)). This behavior is similar to that reported by Lee et al. [13].
@@
In order to characterize this BSW, we systematically measured the WER as a function of current with a magnetic field applied in each of the three directions in space. Fig. 3 shows the WER maps obtained for the same device. The first thing to notice is that the different maps are not exactly symmetrical for positive and negative currents. This is due to the presence of a small positive DC bias current, I µA injected continuously into the device to measure its AHE resistance. However, the disturbance caused is small compared to the switching current ( µA) and does not prevent analysis. The WER measured as a function of an H field for a positive current or a negative current (Fig. 3 (a)) shows 3 regions. The blue region, corresponding to a maximum value of the WER = 1, is the sub-critical region for which the applied current is less than the magnetization reversal critical current in the presence of the applied H field. The yellow region, corresponding to a minimum value of the WER, is the deterministic magnetization reversal region. Finally, the light green-blue region, corresponding to an intermediate value of the WER, is the region where the magnetization begins to switch back towards its initial state. As the H field increases, the energy barrier to be crossed for magnetization switching decreases while the initial amplitude of the DL term increases, leading to a decrease in the critical reversal current. This is observed at the boundary between the blue and yellow regions. In addition, this decrease is almost linear with the increase in the applied magnetic field, as anticipated by macrospin modeling [19, 30, 31]. The behavior of the boundary between the yellow and light green-blue regions is more difficult to determine here. For negative currents, the amplitude of the current for which the BSW appears, the BSW threshold current, decreases with an increase in the applied H field. For positive currents, no clear trend is observed. These three regions can be found on all the maps shown in Fig. 3.
@@
Fig. 3 (b) shows the WER maps as a function of a variable H field and increasing SOT current for an applied H field of 600 Oe. When the applied H field is positive or slightly negative (respectively negative or slightly positive) for a positive (respectively negative) applied current, the reversal of the magnetization is stable and deterministic (yellow region). When this applied field changes direction, the WER increases sharply and no deterministic magnetization switching can be obtained. Moreover, for the deterministic reversal regions, the amplitude of the critical reversal current decreases with that of the applied magnetic field, whereas the opposite behavior is observed for the boundary between the switching and back-switching regions. This behavior is consistent with a negative effective field, H, produced by a positive injected current: when the applied H field adds to H, the energy barrier separating the two magnetization states decreases, leading to a reduction in the critical reversal current [13, 25, 32]. However, as the amplitude of the resulting H field increases, the magnetization approaches an in-plane position at the end of the pulse, resulting in a loss of determinism, as explained in the next section.
@@
Finally, Fig. 3 (c) shows the WER maps as a function of a variable H field and SOT current for an applied H field of 800 Oe. These maps are in agreement with the hypothesis of an up (respectively down) magnetization orientation stabilized by a positive (respectively negative) magnetic field, H. Thus, for positive (respectively negative) currents, the critical reversal current increases (respectively decreases) when the amplitude of the negative magnetic field increases. More (or less) current is needed to stabilize an up (or down) magnetization in the presence of a negative H field. Similar reasoning applies to positive fields. This hypothesis is consistent with the negative sign of the spin Hall effect in W and the presence of a positive H field [33]. We can also see the presence of a blue zone for high positive fields and large negative currents, a zone that also exists for opposite current and H field. During the current pulse, the FL and DL spin orbit torques maintain the magnetization in a dynamic equilibrium whose direction approaches the equator (m=0) as the current amplitude increases (see the section on macrospin simulations below). At the end of the current pulse, only the applied magnetic field, H, remains, dictating the final orientation of the magnetization.

### v3

@@ I Introduction
Magnetic random-access memories (MRAMs) are considered one of the most promising replacements for volatile CMOS memory technologies across all levels of the memory hierarchy. Among the latest generations of MRAMs, spin-orbit torque MRAMs (SOT-MRAMs) are particularly attractive for replacing memories closest to the processing units, as they combine high endurance with sub-nanosecond switching times [prenat_ultra-fast_2016, shao_roadmap_2021]. Nevertheless, while spin transfer torque MRAMs (STT-MRAMs) are commercially available [noauthor_spin-transfer_nodate], SOT-MRAMs must overcome several challenges before they can be adopted by the industry [shao_roadmap_2021]. One of these major challenges is to achieve deterministic switching, meaning reliably controlling the result of the write operation, over a wide current range. The key metric for deterministic switching is the write error rate, which measures the probability of getting the expected final state after a write operation.
@@
In SOT-MRAM technology with perpendicular magnetic anisotropy, which is the most promising solution in terms of scaling, density and data retention, magnetization switching by means of spin-orbit current alone is not deterministic. Indeed, by symmetry, the probabilities of obtaining an ”up” or ”down” magnetization are equal. An additional factor to break this symmetry is required. In early implementations, this breaking of symmetry was induced by the application of an in-plane magnetic field, collinear with the current but external to the system [miron_perpendicular_2011, cubukcu_spin-orbit_2014, liu_spin-torque_2012]. As this solution is not suitable for embedded applications, a number of propositions have been published in recent years, aimed at developing deterministic and integrated ”field-free” SOT-MRAM cells, based on tilted anisotropy [you_switching_2015], combination of SOT and STT effects [cai_sub-ns_2021], exchange bias [van_den_brink_field-free_2016], embedded magnet in the material stack [garello_manufacturable_2019] and unconventional torques from new materials [kao_deterministic_2022] (for a more complete list, please consult review articles such as [krizakova_spin-orbit_2022]).
@@
However, the presence of symmetry breaking does not guarantee a perfectly deterministic reversal of magnetization, and consequently a deterministic writing. It has been observed that increasing the write current beyond the threshold for reversal leads, as expected, to an increase in the probability of magnetization switching. However, if the current is increased further, the probability then decreases and may even oscillate as a function of the applied current [lee_oscillatory_2018, yoon_anomalous_2017, decker_time_2017]. Similar back-switching (BSW) phenomenon has previously been observed and studied in STT-MRAMs [min_back-hopping_2009]. One origin of this phenomenon has been attributed to the instability of the reference layer magnetization in magnetic tunnel junctions (MTJ), an instability that can even lead to the reversal of this magnetization [kim_experimental_2016, devolder_back_2020]. In SOT-MRAMs, magnetization is switched by means of a current injected in the plane of the layers [miron_perpendicular_2011, liu_spin-torque_2012], without the need for an MTJ-type structure to obtain a spin-polarized current. Observation of the back-switching phenomenon in simple structures, such as non-magnetic/ferromagnetic/insulator trilayer stacks, demonstrates that in SOT-MRAM this BSW is intrinsic to the magnetization dynamics induced by spin-orbit torques and not related to the dynamics of the reference layer. This BSW phenomenon is a challenge for the efficient design of memories based on SOT-MRAMs (Fig. 1 (a)). Its study is essential to our understanding of the magnetization dynamics induced by spin-orbit torques, and for finding effective means of mitigating BSW in SOT-MRAMs.
@@
Despite its importance, relatively few studies have been devoted to it. Most are numerical studies based on macrospin simulations. The existence of this BSW phenomenon was reported as early as the first macrospin SOT switching simulations, whether with or without the Field-Like (FL) term [lee_threshold_2013, park_macrospin_2014]. The importance and the role of the FL term on the BSW phenomenon were then highlighted [legrand_coherent_2015, taniguchi_theoretical_2019, zhu_threshold_2020]. The existence of this BSW phenomenon has been explained by energy arguments: at the end of the write current pulse, if the energy of the system is higher than the saddle point energy, the magnetization will relax in either a first potential well corresponding to a first direction of magnetization or in the other corresponding to the opposite state of magnetization [park_macrospin_2014, taniguchi_theoretical_2019]. This explains the high sensitivity of the phenomenon to damping, as reported in particular by [lee_threshold_2013, park_macrospin_2014, taniguchi_reduction_2020]. Experimentally, this phenomenon has been reported in some publications - for example, in Pt/Co/AlO layers, where the magnetization is reported to switch back to its original orientation above a certain threshold of applied planar magnetic field [decker_time_2017]; or in Ta/CoFeB/MgO-based magnetic tunnel junctions, where a decrease in the probability of the magnetization switching is observed when the write current is increased [krizakova_tailoring_2022]. Only two publications from the same group report a more complete study of this phenomenon, combining experiments and simulations [yoon_anomalous_2017, lee_oscillatory_2018]. Both studies were carried out on micron-sized devices patterned from a Ta/CoFeB/MgO based stack. Time-resolved MOKE measurements show a decrease in the probability of switching as the writing current increases. Using analytical modeling and micromagnetic simulations, these measurements were explained by a BSW phenomenon due to the reflection of the magnetic domain walls (DW) on the edges of the device. This reflection can only occur for a negative sign of the ratio of the FL and Damping-Like (DL) torques [yoon_anomalous_2017]. In a subsequent study, based on electrical measurements on Hall crosses supporting a magnetic pillar, the authors observed an oscillatory and deterministic behavior of the magnetization switching while increasing the amplitude of nanosecond write pulses [lee_oscillatory_2018].
@@
In this work, we present a systematic study of the back-switching phenomenon using a statistical approach. The measurements were carried out on sub-100nm CoFeB magnetic pillars at the center of -W Hall crosses. The error rate of the write operation was measured as a function of various parameters such as the amplitude of the write current pulse and an external magnetic field applied along different directions. We show that macrospin simulations, using realistic values for the different magnetic parameters, reproduce the experimental observations well. These macrospin simulations open the door to the development of compact models that take this BSW phenomenon into account. These numerical simulations give us an insight into the mechanism behind the BSW phenomenon. They also allow us to study ways of reducing this phenomenon, such as using stronger damping, tuning the FL/DL ratio or, following previous work from [zhu_threshold_2020] controlling the shape of the writing pulses. We test the latter solution and show experimentally that this strategy does indeed mitigate back-switching and increases the current range for write operations in SOT-MRAMs.
@@ II Methodology
The measurements were carried out on a stack constituting the typical storage layer of a SOT-MRAM: -W(4)/CoFeB(0.9)/MgO(RA20 m2)/CoFeB(0.3)/ Ta(1)/Ru(5) where the thicknesses are expressed in nanometers. The top 0.3 nm CoFeB layer is non-magnetic and is used to provide appropriate environment to the MgO layer during annealing. The layers were sputter deposited and annealed at 350° C for 30 minutes unless otherwise specified in the text. This stack was then patterned as a CoFeB/MgO/CoFeB/Ta/Ru pillar at the center of a -W Hall cross by successive steps of electron lithography and ion beam etching. The studies reported here, except for certain cases specified in the text, were carried out on pillars with a real diameter of around 60 nm sitting on a 100 nm wide track. Fig. 1 (b) presents the magnetic stack, a representative SEM image of a studied device, and a schematic of the measurement setup.
The study is based on statistical measurements of Write Error Rate (WER), which is a crucial metric for qualifying memory devices. Typically, while this WER is very low in SRAM memories, it is much higher in MRAMs. However, to operate as a working memory, the WER of a single bit should typically be less than 10-9 if the chip incorporates error correction code (ECC), or less than 10-18 without this error correction [khvalkovskiy_basic_2013]. Fig. 2 (a) shows the WER measurement scheme that has been used and its definition: the memory point is initialized to a defined state and this state is read to verify that initialization was successful. A pulse to write the opposite state, of variable amplitude, is then applied and the state of the memory point is read again. WER is defined as the ratio of the number of write errors after a successful initialization to the number of successful initializations. Note that a WER of 0.5 corresponds to an equal probability of obtaining the two states of opposite magnetization.
@@ III Experimental Results
Magnetic characterization of the samples confirms perpendicular anisotropy, as evidenced by square hysteresis loops, shown in [supplementary]. The coercive field is approximately 750 Oe, while the anisotropy field (H) and the thermal stability factor (), extracted using the switching field distribution method [tillie_data_2016, feng_sweep-rate-dependent_2004], are estimated to be around H=2600 Oe and =45, respectively [supplementary]. A typical hysteresis curve for current induced magnetization reversal is shown in Fig. 2 (b). The areas colored in green represent known areas of magnetization reversal into the opposite state. However, as the amplitude of the applied voltage increases, a new behavior appears, represented by the areas colored in red. The final state of magnetization is no longer fixed, opposite to the initial state, but begins to fluctuate in an apparently random manner between these two states. This fluctuation occurs before eventually switching back to the initial state, as observed for positive voltages. This behavior is not observed for negative voltages, likely due to a slight asymmetry in the experimental setup, as discussed below.
@@
The corresponding WER measurement is shown in Fig. 2 (c) for positive voltages. This curve was obtained with 100 write/read cycles, which explains the saturation at WER, representing in this case an absence of error. Starting from a down state and applying a positive voltage, the WER of the up state is measured. We first observe a zone without switching when the voltage is below a first threshold V0. Then when the voltage is increased above V0, we observe that, after a transition zone, the magnetization switches to the up state without error, corresponding to the green zone in Fig. 2 (b). When the voltage increases further, there is another threshold V1 at which the error rate starts increasing and approaches the maximum value 1, which corresponds to the magnetization switching back to its initial orientation, in this case down, without any error. This behavior is repeated when the magnetization is initialized in the opposite up state and by measuring the WER of the down state. By applying a positive voltage below V1, the up state remains stable while above V1, the magnetization switches to the down state. This shows that the BSW is not random but, on the contrary, deterministic: after being reversed to the orientation opposing the initial state, the magnetization can be switched back to its initial orientation in a deterministic way, as shown by the low error rate obtained (cf Fig. 2 (c)). This behavior is similar to that reported by Lee et al. [lee_oscillatory_2018].
@@
In order to characterize this BSW, we systematically measured the WER as a function of current with a magnetic field applied in each of the three directions in space. Fig. 3 shows the WER maps obtained for the same device. The first thing to notice is that the different maps are not exactly symmetrical for positive and negative currents. This is due to the presence of a small positive DC bias current, I µA injected continuously into the device to measure its AHE resistance. However, the disturbance caused is small compared to the switching current ( µA) and does not prevent analysis. The WER measured as a function of an H field for a positive current or a negative current (Fig. 3 (a)) shows 3 regions. The blue region, corresponding to a maximum value of the WER = 1, is the sub-critical region for which the applied current is less than the magnetization reversal critical current in the presence of the applied H field. The yellow region, corresponding to a minimum value of the WER, is the deterministic magnetization reversal region. Finally, the light green-blue region, corresponding to an intermediate value of the WER, is the region where the magnetization begins to switch back towards its initial state. As the H field increases, the energy barrier to be crossed for magnetization switching decreases while the initial amplitude of the DL term increases, leading to a decrease in the critical reversal current. This is observed at the boundary between the blue and yellow regions. In addition, this decrease is almost linear with the increase in the applied magnetic field, as anticipated by macrospin modeling [lee_threshold_2013, taniguchi_critical_2015, torrejon_current-driven_2015]. The behavior of the boundary between the yellow and light green-blue regions is more difficult to determine here. For negative currents, the amplitude of the current for which the BSW appears, the BSW threshold current, decreases with an increase in the applied H field. For positive currents, no clear trend is observed. These three regions can be found on all the maps shown in Fig. 3.
@@
Fig. 3 (b) shows the WER maps as a function of a variable H field and increasing SOT current for an applied H field of 600 Oe. When the applied H field is positive or slightly negative (respectively negative or slightly positive) for a positive (respectively negative) applied current, the reversal of the magnetization is stable and deterministic (yellow region). When this applied field changes direction, the WER increases sharply and no deterministic magnetization switching can be obtained. Moreover, for the deterministic reversal regions, the amplitude of the critical reversal current decreases with that of the applied magnetic field, whereas the opposite behavior is observed for the boundary between the switching and back-switching regions. This behavior is consistent with a negative effective field, H, produced by a positive injected current: when the applied H field adds to H, the energy barrier separating the two magnetization states decreases, leading to a reduction in the critical reversal current [lee_oscillatory_2018, krizakova_tailoring_2022, fan_asymmetric_2019]. However, as the amplitude of the resulting H field increases, the magnetization approaches an in-plane position at the end of the pulse, resulting in a loss of determinism, as explained in the next section.
@@
Finally, Fig. 3 (c) shows the WER maps as a function of a variable H field and SOT current for an applied H field of 800 Oe. These maps are in agreement with the hypothesis of an up (respectively down) magnetization orientation stabilized by a positive (respectively negative) magnetic field, H. Thus, for positive (respectively negative) currents, the critical reversal current increases (respectively decreases) when the amplitude of the negative magnetic field increases. More (or less) current is needed to stabilize an up (or down) magnetization in the presence of a negative H field. Similar reasoning applies to positive fields. This hypothesis is consistent with the negative sign of the spin Hall effect in W and the presence of a positive H field [hao_giant_2015]. We can also see the presence of a blue zone for high positive fields and large negative currents, a zone that also exists for opposite current and H field. During the current pulse, the FL and DL spin orbit torques maintain the magnetization in a dynamic equilibrium whose direction approaches the equator (m=0) as the current amplitude increases (see the section on macrospin simulations below). At the end of the current pulse, only the applied magnetic field, H, remains, dictating the final orientation of the magnetization.